\section*{References}
\begin{thebibliography}{99}
\bibitem{PL}  H. J. de Vega and L. N. Lipatov, 
Phys. Lett. {\bf B578}, 335 (2004).
 
 \bibitem{lipa} L. N. Lipatov, Sov. J. Nucl. Phys. ${\bf 23}$ (1976)
  642; V. S. Fadin, E. A. Kuraev and L. N. Lipatov, Phys. Lett. {\bf B60}, 50 (1975);
Sov. Phys. JETP {\bf 44} (1976) 443; {\bf 45} (1977) 199;
Ya. Ya. Balitsky and L. N. Lipatov, Sov. J. Nucl. Phys. {\bf 28}, 822 (1978).
 L. N. Lipatov, Phys. Lett. {\bf B251}, 284 (1990).
 L. N. Lipatov, Phys. Rep., {\bf 286} (1997) 132.
 L. N. Lipatov, Sov. Phys. JETP {\bf 63} (1986) 904.
\bibitem{bh} See for example,
N. D. Birrell, P. C. W. Davis, `Quantum Fields in Curved Space',
Cambridge U. Press, 1982 and references therein.
S. W. Hawking, Comm. Math. Phys. 43, 199 (1975).
N. Sanchez, Phys. Lett. B87, 212 (1979), Phys. Rev. D24, 2100 (1981).

\bibitem{QGP}  {\it Quark-Gluon Plasma 3}, edited by R. C. Hwa and X. N. Wang,
World Scientific. E. V. Shuryak, World Sci. Lect. Notes Phys. 8: 1-401, 1988.
H. Satz, Rept. Prog. Phys. 63, 1511 (2000). E. V. Shuryak,
I. Zahed, Phys. Rev. C70, 021901 (2004) and  D. Boyanovsky,  H. J. de Vega,
Phys. Rev. {\bf D68}, 065018  (2003) and Nucl. Phys. A747, 564 (2005).

\bibitem{BKP}  J. Bartels, Nucl. Phys. ${\bf {B175}}$ (1980) 365;
J. Kwiecinski and M. Prascalowicz, Phys. Lett. ${\bf {B94}}$ (1980) 413.

\bibitem{integr}  L. N. Lipatov, hep-th/9311037, Padova preprint
DFPD/93/TH/70, unpublished.

\bibitem{Heis}  L. N. Lipatov, Sov. Phys. JETP Lett. ${\bf {59}}$ (1994) 571.
 L. D. Faddeev, G. P. Korchemsky, Phys. Lett. {\bf {342}} (1995) 311.

\bibitem{lip85}  L. N. Lipatov, Sov. Phys. JETP, 63, 904 (1986) [Zh. ETF, 90, 1536 (1986)].

\end{thebibliography}

\documentclass[11pt]{article}
\usepackage{blois,epsfig}

\def\be{\begin{equation}}
\def\ee{\end{equation}}
\def\bea{\begin{eqnarray}}
\def\eea{\end{eqnarray}}

\begin{document}

\vspace*{2cm}
\title{QCD Pomerons at Non-Zero Temperature}

\author{H. J. de Vega$^a$ and L. N. Lipatov$^b$}

\address{$^a$ Laboratoire de Physique Th\'eorique et Hautes 
Energies, Universit\'es Paris VI et VII, BP 126,\\
4, place Jussieu, F-75252 Paris cedex 05, France.\\
Laboratoire Associ\'e au CNRS UMR 7589. \\
$^b$ Theoretical Physics Department,
Petersburg Nuclear Physics Institute,\\
Gatchina, 188300, St. Petersburg, Russia}

\maketitle\abstracts{We consider the QCD scattering amplitudes at high energies $\sqrt{s}$
and fixed momentum transfers $\sqrt{-t}$ in the leading logarithmic
approximation at a non-zero temperature $T$ in the $t$-channel.  It is
shown that the BFKL Hamiltonian has the property of holomorphic
separability. The holomorphic Hamiltonian for $n$-reggeized gluons at 
temperature $T$ is shown to  be obtained from the $T=0$ Hamiltonian by
an unitary transformation. The connection with field quantization
in accelerated frames and black hole backgrounds is discussed.
The effect of the running of $\alpha_{QCD}$ on the pomeron eigenvalues is investigated.}

\section{Quark Gluon Plasma and Reggeons}

Current theoretical understanding of the quark-gluon plasma (QGP)
generated in heavy nucleus collisions
suggests that the QGP thermalizes via parton-parton
scattering. The QGP is understood to  cool down by hydrodynamic
expansion till the temperature reaches the hadronization scale $\sim 160$MeV.
One interesting phenomena is the suppression of the $\psi $-meson
production in the heavy nucleus collisions due to the
disappearance of the confining potential between $q$ and $\bar{q}$ at
high temperature \cite{QGP}. 
A similar effect should exist for glueballs
constructed from gluons. Because the Pomeron is considered as a composite
state of reggeized gluons, the influence of the temperature on its
properties is of great interest. We constructed in  ref.\cite{PL} the BFKL
equation at temperature $T$ in the center-mass system of the $t$
-channel (where $\sqrt{t}=2\epsilon $) and show that the BFKL dynamics in a 
thermostat for composite states of $n$ reggeized gluons in multi-colour QCD
is integrable. 

Let us consider the Regge kinematics in which the total particle
energy $\sqrt{s}$ is asymptotically large in comparison with the
temperature $T$. In this case one can neglect
the temperature effects in the propagators of the initial and intermediate
particles in the direct channels $s$ and $u$. But the momentum transfer
$|q|$ is considered to be of the order of $T$ (note, that $q_{\mu }$
is the vector orthogonal to the initial momenta $q_{\mu }\approx
q_{\mu }^{\perp }$ ). As it is well known the particle wave
functions $\psi (x_{\mu })$ at temperature $T$ are periodic in the
euclidean time $\tau =i\,t$ with period $1/T$.

We introduce the temperature $T$ in the center of mass frame of the
$t$-channel. Thus, the euclidean energies of the intermediate
gluons in the $t$-channel become quantized as
$$
k_{4}^{(l)}=2\pi l\,T \; .
$$
In the $s$-channel the
invariant $t$ is negative and therefore the analytically continued 4-momenta
of the $t$-channel particles can be considered as euclidean vectors. It
means, that at temperature $T$, the wave functions
for virtual gluons are periodic functions of the 
holomorphic impact-parameter
$\rho =x+iy$ 
with imaginary period $\frac{i}{T}$. Also,
the canonically conjugated momenta $p$ have their imaginary part quantized,
\begin{equation}
\rho \rightarrow \rho +\frac{i}{T} \quad , \quad p=\mbox{Re}\, p+ 
\pi\, i \, l \, T \; .
\end{equation}
with integer $l$ (note that $p=(p_1+ip_2)/2 $).

\section{Reggeon Hamiltonian at temperature $T$}

The calculation of the Regge trajectory $1+\omega (t)$ of the gluon 
at temperature $T$ in the t-channel,
in one-loop approximation reduces to the integration over the
real part $k_{1}$ of the transverse momentum $ k $
of the virtual gluon and to the summation over its imaginary part
$k_{2}=l$. In such a way we obtain
the following result for the trajectory having the separability
property [cf. \cite{lipa}],
\begin{equation}
\omega (-\vec{q}^{2})=-\frac{g^{2}}{8\pi ^{2}} \; N_{c} \; \Omega (-\vec{q}
^{2})\quad , \quad \Omega (-\vec{q}^{2})=\Omega (q)+\Omega (q^{\ast }) \; .
\end{equation}
Here,
\begin{equation}
\Omega (q)=\frac{\pi T}{\lambda } + \frac{1}{2}\,\left[ \psi (1+\frac{iq}{2\pi\,T})
+\psi(1-\frac{iq}{2\pi\,T})-2\psi (1)\right] \; , \nonumber
\end{equation}
where we regularized the infrared divergence for the zero mode $l=0$
introducing a mass $\lambda $ for the gluon (see \cite{lipa}).

A similar  divergence appears in the Fourier transformation $G(
\overrightarrow{\rho }_{12})$ of the effective gluon propagator $(\vec{k}
_{\perp }^{2}+\lambda ^{2})^{-1}$ contained in the product of the effective
vertices $q_{1}k^{-1}q_{2}^{\ast }$ for the production of a gluon
with momentum $k_{\mu }$ (cf. \cite{lipa})
\begin{equation}
G(\overrightarrow{\rho }_{12})=-\frac{\pi \,T}{\lambda }+\ln \left( 2\sinh
\pi \, T \, \rho _{12}\right) +\ln \left( 2\sinh \pi \, T \, \rho _{12}^{\ast }
\right) \; .
\end{equation}
Therefore, the divergence at $\lambda \rightarrow 0$ cancels in the
sum of kinetic and potential contributions to the BFKL equation and the
Hamiltonian $H_{12}$ for the Pomeron in a thermostat has the property of
holomorphic separability with the holomorphic Hamiltonian given below 
(as it is the case for zero temperature \cite{lipa})
\begin{equation}
h_{12}=\sum_{r=1}^{2}\left[ \frac12 \; \psi (1+\frac{ip_{r}}{2\pi\,T})+
\frac12 \; \psi(1-\frac{ip_{r}}{2\pi\,T}) + %\right. \cr \cr &+&\left. 
\frac{1}{p_{r}}\,\ln \left( 2\sinh \pi \, T \,\rho_{12}\right) \,p_{r}-
\psi(1)\right] \; .  \label{hamilT}
\end{equation}
We constructed in ref. \cite{PL} the eigenfunctions and eigenvalues of
the hamiltonian eq. (\ref{hamilT}) in terms of hypergeometric functions.

\section{Wavefunctions in coordinate space}

The Pomeron wave function can be constructed directly in coordinate
space. For this purpose we use the conformal transformation
\begin{equation}\label{exp}
\rho_{r}^{\prime }= \frac1{2 \, \pi \, T} \; e^{2 \, \pi \, T \, \rho _{r} }
\end{equation}
to map the zero temperature wave functions into the wavefunctions at temperature
$T$.

Thus, the Pomeron wave function at non-zero
temperature having the property of single-valuedness and periodicity
takes the form
\begin{equation}
\Psi ^{(m,\widetilde{m})}(\vec{\rho _{1}},\,\vec{\rho _{2}},\,\vec{\rho_{0}})= 
\left( \frac{\sinh [\pi \, T \, \rho_{12}]}{2\sinh [\pi \, T \, \rho _{10}]
\,\sinh [\pi \, T \, \rho _{20}]}\right) ^{m} \times 
\left( \frac{\sinh [\pi \, T \, \rho
_{12}^{\ast }]}{2\sinh [\pi \, T \, \rho _{10}^{\ast }] \,\sinh [\pi \, T \, \rho
_{20}^{\ast }]}\right) ^{\widetilde{m}}\,.
\end{equation}
The orthogonality and completeness relations for these functions can be
easily obtained from the analogous results for $T=0$ (see \cite{lipa})
using the above conformal transformation. 

Moreover, the pair BFKL Hamiltonian $h_{12}$ can be expressed in terms of
the BFKL Hamiltonian at zero temperature in the new variables
\begin{equation}
h_{12}=   \ln (p_{1}^{\prime }\,p_{2}^{\prime })+   %      \\ \cr &&+
\frac{1}{p_{1}^{\prime }}
\,\log (\rho _{12}^{\prime })\,\,p_{1}^{\prime }+\frac{1}{p_{2}^{\prime }}
\,\log (\rho _{12}^{\prime })\,\,p_{2}^{\prime }-2\psi (1)\,, \nonumber
\end{equation}
where $p_{r}^{\prime }=i\frac{\partial }{\partial \rho _{r}^{\prime
}}$. We recognize here the zero temperature BFKL Hamiltonian \cite{lipa}.
In the course of the derivation the 
following operator identity (see \cite{lipa})
\[
\frac{1}{2}\left[ \psi \left(1+z\frac{\partial }{\partial z}\right)+\psi
  \left(-z\frac{ \partial }{\partial z}\right)\right] =\ln \,z+\ln
  \,\frac{\partial }{\partial z}
\]
was used to transform the kinetic part as well as  properties of the
$\psi $-function.

In summary, the exponential mapping eq.(\ref{exp}) 
maps the reggeon dynamics from zero temperature to temperature
$T$. This mapping explicitly exhibits a periodicity 
$\rho \rightarrow \rho +\frac{i}{T}$ for a thermal state. 
It must be noticed that such class
of mappings are known to describe thermal situations for
quantum fields in accelerated frames and in black hole backgrounds\cite{bh}.

\section{Integrability at non-zero temperature}

As it is well known\cite{BKP}, the BFKL equation at $T=0$ can be
generalized to composite states of $n$ reggeized gluons. In
the multi-colour limit $N_{c}\rightarrow \infty $ the BKP equations are
significantly simplified thanks to their conformal invariance,
holomorphic separability and integrals of
motion \cite{lipa}. The generating function for the holomorphic
integrals of motion
coincides with the transfer matrix for an integrable lattice spin model
\cite{integr}. The transfer matrix is the trace of the
monodromy matrix
\[
t(u)=L_{1}(u)\,L_{2}(u)...L_{n}(u)\,,
\]
satisfying the Yang-Baxter equations \cite{integr}. The integrability of the
$n$-reggeon dynamics in multi-colour QCD is valid also at non-zero
temperature $T$, where, according to the above arguments we should
take the $L$-operator in the form
\[
L_{k}=\left(
\begin{array}{cc}
u+p_{k} & e^{-2 \, \pi \, T \, \rho _{k}}\,p_{k} \\
-e^{2 \, \pi \, T \, \rho _{k}}\,p_{k} & u-p_{k}
\end{array}
\right) \,.
\]
In particular, the holomorphic Hamiltonian is the local Hamiltonian of the
integrable Heisenberg model with the spins being unitarily transformed
generators of the M\"{o}bius group (cf. \cite{Heis})
\begin{eqnarray}
&& M_{k}=\partial _{k} \quad ,  \\ \cr
&& M_{+}=e^{-2 \, \pi \, T \, \rho _{k}}\,\partial
_{k} \quad ,  \quad M_{-}=-e^{2 \, \pi \, T \, \rho _{k}}\,\partial _{k}\; .
\nonumber
\end{eqnarray}
Because the Hamiltonian at  non-zero temperature can be obtained by an
unitary transformation from the zero temperature Hamiltonian, the
spectrum of the intercepts
for multi-gluon states is the same as for zero temperature 
and the wave functions of the composite states can be calculated
by the substitution $\rho _{k}\rightarrow \frac1{2 \, \pi \, T} \;
e^{2 \, \pi \, T \; \rho _{k}}$. 

\section{Running with the QCD coupling}
Taking into account the  running of $\alpha_{QCD}$ 
to one-loop level $\alpha_s (Q) = \frac{4 \, \pi}{9 \, \log \frac{4 \, Q^2}{\Lambda^2}} $
changes the previous results. Now, the pomeron wave function must be an eigenfunction of the operator\cite{lip85}
\be\label{ope}
{\cal E} (Q) \equiv \alpha_s (Q) \; \chi\left(- \frac{i}2 
\frac{d}{d \ln |\rho| }\right) \quad , \quad \chi(\nu)= -\frac6{\pi} {\rm Re}\left[\gamma+\psi\left(\frac12+i \, \nu\right)
\right]
\ee
Semiclassically, we have as quantization condition 
\be
\phi(\nu_k) = k + \frac14 \quad , \quad k=0,1,2,\ldots
\ee
where
\be
\phi(\nu) \equiv \frac4{9 \, \alpha_s (Q)}\left[ \frac1{\chi(\nu)} \;
\int_0^{\nu} dx \; \chi(x) - \nu\right] - \frac1{2 \, \pi} \;
\delta_{\nu}^T(\vec{Q}) 
\ee
and
$$
 e^{i \, \delta_{m, {\tilde m}}^T(\vec{Q})} = |Q|^{-4 \, i \, \nu} \;
 \frac{\Gamma(m+i \, Q)}{\Gamma(1-m+i \, Q)}
\frac{\Gamma( {\tilde m}-i \, Q^*)}{\Gamma(1- {\tilde m}-i \, Q^*)}   \; .
$$
The corresponding eigenvalue of the operator $ {\cal E} (Q) $ are given
by
\be
\omega_k(Q) =  \alpha_s(Q) \; \chi(\nu_k)
\ee
The eigenvalues and eigenfunctions now depend parametrically on $ Q 
= \frac{Q_{phys}}{2 \, \pi \, T} $. We recover the zero temperature
limit for $ Q \to \infty $. 
The characteristic scale in temperature is given by $
T_{chara} \equiv \frac1{4 \, \pi} \, e^{\frac{2 \, \pi}{9 \, \alpha_s(Q)}} $
We find that the eigenvalue for $k=0$  first grows
with the temperature and then goes down. In addition, we see that rotational 
invariance is recovered for $ T \to 0 $ since the results for Im Q $= 0 $ and 
$1$ coincide in such limit.

\end{document}